OpenStudy (kramse18):

The acceleration function for a particle moving along a line is a(t)=2t+1 . The initial velocity is v(0)=−12. Then: The velocity at time t, v(t)= . The distance traveled during the time interval [0,5] is equal to =

OpenStudy (noelgreco):

Integrate twice.

OpenStudy (kramse18):

which I did and I'm not getting the right answer

OpenStudy (noelgreco):

After the first integration, what did you get for v?

OpenStudy (kramse18):

t^2+t-12

OpenStudy (noelgreco):

good. Now integrate that with integration limits as provided. What was your answer?

OpenStudy (kramse18):

(1/3)t^3+(1/2)t^2-12t

OpenStudy (noelgreco):

Good. I got 35/6

OpenStudy (noelgreco):

Sorry, \[\frac{ -35 }{ 6 }\]

OpenStudy (kramse18):

says it isn't right because that is what I had gotten first off

OpenStudy (noelgreco):

Is your answer from a book or an instructor?

OpenStudy (kramse18):

its online homework

OpenStudy (astrophysics):

\[v(t) = \int\limits 2t+1 dt = t^2+t+C \] Note the initial conditions are v(0)=-12 \[\implies v(t)=t^2+t-12\] that will be your velocity, now you need to find distance, again integrating through \[x(t) = \frac{ t^3 }{ 3 }+\frac{ t^2 }{ 2 }-12t \] we have to note the change of direction when v = 0, that's the trick here.

OpenStudy (noelgreco):

Astro is right. The particle changes direction when v=0. Negative areas count in the distance moved. I may be getting rusty.

OpenStudy (kramse18):

so then what do you do for that?

OpenStudy (noelgreco):

First, solve for v(t) = 0

OpenStudy (astrophysics):

The wording is a bit odd, they should write displacement, so you can find all the individual displacements and add them up

OpenStudy (noelgreco):

Got to go to dinner - will check back, but it looks like Astro has the key.

OpenStudy (kramse18):

Astrophysics I need some help

OpenStudy (astrophysics):

First, when does the particle change direction?

OpenStudy (kramse18):

at 0

OpenStudy (astrophysics):

Right v(t) = 0 so what do you get

OpenStudy (kramse18):

-12?

OpenStudy (astrophysics):

I meant \[v(t) = 0 \implies 0 = t^2+t-12\]

OpenStudy (astrophysics):

What do you get

OpenStudy (kramse18):

t=3 and t=-4

OpenStudy (astrophysics):

So notice -4 is not in our time interval, so now we'll have two integral from 0 to 3 and 3 to 5

OpenStudy (astrophysics):

Or you can find individual displacements and add them up

OpenStudy (astrophysics):

So find x(3)-x(1), x(5)-x(3), and x(1) and add them up

OpenStudy (astrophysics):

That will give you your total distance

OpenStudy (kramse18):

do you put those numbers in the first or second anti-derivative?

OpenStudy (astrophysics):

You want the distance, so distance

OpenStudy (kramse18):

i'm still getting the wrong answer. Sorry I'm not trying to be complicated, I just don't fully understand

OpenStudy (astrophysics):

Can you show me your work

OpenStudy (kramse18):

Well I plugged the (1/3)t^3+(1/2)t^2-12t into the calculator to get those points you said. 5=-5.8333 3=-22.5 1=-11.17, subtracted the 2 intervals you said and then added together what I got from that

OpenStudy (astrophysics):

Ok one second, so we have v(t) = (t-3)(t+4) correct

OpenStudy (astrophysics):

Lets do \[0<t<1, ~~~~ 1<t<3,~~~~ 3<t<5\] we can find the individual displacements here

OpenStudy (astrophysics):

\[x(1)-x(0) = -67/6 \] \[x(3)-x(1) = -45/2-(-67/6) = -34/3\] \[x(5)-x(3) = 16/3\] check the math

OpenStudy (astrophysics):

Your total distance is the absolute value of all that

OpenStudy (astrophysics):

Hey can you check your initial condition

OpenStudy (kramse18):

condition? [0,5]

OpenStudy (astrophysics):

v(0)

OpenStudy (astrophysics):

How many attempts do you have btw?

OpenStudy (kramse18):

its unlimited attempts

OpenStudy (astrophysics):

Ok can you try 235/6

OpenStudy (kramse18):

that was right

OpenStudy (astrophysics):

Ok perfect, lets just stick to my first though then, we'll integrate through and get \[- \int\limits_{0}^{3} t^2+t-12 dt - \int\limits_{3}^{5} t^2+t-12dt \]

OpenStudy (astrophysics):

Because the other way might be very confusing especially since i'm not doing it on paper

OpenStudy (astrophysics):

\[- \int\limits\limits_{0}^{3} t^2+t-12 dt \color{red}{+} \int\limits\limits_{3}^{5} t^2+t-12dt \]

OpenStudy (astrophysics):

That's your distance

OpenStudy (kramse18):

how did you get those intervals?

OpenStudy (astrophysics):

Good question, alright so remember when I asked you, to find where the particle changes direction, we need to set v(t) = 0 that gave us 0=(t-3)(t+4) which gives 3, -4 we can neglect the -4 since it is not in our interval [0,5], but the 3 here tells us when the particle changes direction you can even try plugging in the values in v(t) to check your self, you will see 0<t<3 we get v<0, then 3<t<5 we get v>0 and we have that negative in front of the first integral because other wise we will get a negative value, but here we want the distance so we are taking the positive values or essentially the absolute value, you can note \[\int\limits_{a}^{b} |v(t)|dt = distance\]and \[\int\limits_{a}^{b} v(t) dt = displacement\]

OpenStudy (kramse18):

ok thanks. I understand now

OpenStudy (astrophysics):

Yeah haha sorry about all the mess earlier, should've just went with my first though

OpenStudy (astrophysics):

thought*